OpenStudy (anonymous):

I REALLY NEED HELP PLEASE HELP! Mahesh wants to inflate twenty-five foil balloons with helium. If each balloon requires zero point 4 5 cubic feet of helium, how much total helium will he use?

OpenStudy (ghazi):

total amount of helium used = 0.25*25=6.25 cubic feet

OpenStudy (ghazi):

@twilaswift clear?

OpenStudy (anonymous):

one balloon requires 0.45 cubic feet of helium, so two balloons would need twice this amount of gas ( 2 x 0.45 cubic feet), hence by the same reasoning 25 balloons would need 25 x 0.45 cubic feet of helium.

OpenStudy (anonymous):

zero point 1 1 2 5 cubic feet 1 point 1 2 5 cubic feet eleven point 2 5 cubic feet one hundred twelve point 5 cubic feet

OpenStudy (ghazi):

i mean 0.45*25=11.25 cubic feet